OpenStudy (anonymous):

What does "true" mean? For example, we can say "A is true". What does that mean, with regards to A? If A is true, is A true because it is A, or for some other reason?

OpenStudy (anonymous):

5=5 is true or 0=0

OpenStudy (anonymous):

this will help you :) http://www.wolframalpha.com

OpenStudy (anonymous):

did i win the best response

OpenStudy (anonymous):

No, I know that those statements are true, but what does it mean that they are true?

OpenStudy (anonymous):

the = is asking you if it is true or false true means that is right 1/4=.25 is true but 1/4+.25=.5 true or 1/4-.25=0 true

OpenStudy (anonymous):

did that help

OpenStudy (helder_edwin):

that's actully a very interesting question. i don't think it has a conclusive answer. as far i know, it has to do with what one "knows".

OpenStudy (helder_edwin):

in binary logic (aristotelean logic) propositions can be either TRUE or FALSE but not both.

OpenStudy (anonymous):

yes that is true !

OpenStudy (helder_edwin):

try this book (it very cool): "Logic" a very short introduction, by Graham Priest, Oxford University Press, 2000.

OpenStudy (helder_edwin):

also don't forget that recent neuro-biological research has proven that our brain actually carries some "coding" (kind of a computer program) since inception.

OpenStudy (helder_edwin):

so some "truths" might already be programmed into us since our very first day in THIS world.

OpenStudy (helder_edwin):

helpful?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

all thes people and no medal realy

OpenStudy (anonymous):

intredasting...

OpenStudy (radar):

Truth is reality. Untruth is everything else.

OpenStudy (helder_edwin):

yes, to a certain degree. but what about euclidean geometry?

OpenStudy (radar):

Hey I was just trying to say something very profound.

OpenStudy (anonymous):

then give a medal lol

OpenStudy (radar):

You know a long time ago, those guys Plato, Aristole and the rest of that think tank discussed what was real, truth, euclidean geometry, etc., but today we don't have time for that.

OpenStudy (helder_edwin):

euclidean geometry has nothing to do with reality as it IS. it has to do with how our brain works, how it compiles and processes information

OpenStudy (helder_edwin):

i a way i agree with you @radar .

OpenStudy (radar):

That seems to be very internal, and we have hard time getting there, that is why we have specialist who can't analyze themselves, but have a license to analyze others.

OpenStudy (helder_edwin):

the very last part is great.

OpenStudy (radar):

Well where I am it is time to go to bed. Good night all, maybe I can dream the meaning of truth.

OpenStudy (helder_edwin):

cool @radar . though it might be an unsettling dream :)

OpenStudy (radar):

Usually, after awakening, the dreams just evaporate and I can't remember a thing.

OpenStudy (helder_edwin):

have pleasant dreams anyway

OpenStudy (radar):

You too, good night all.